OpenStudy (anonymous):

Megan and Julie are stuck simplifying radical expressions. Megan has to simplify the quantity of x to the one third power, over x to the one twelfth power. Julie has to simplify the thirty second root of the quantity of x times x to the second times x to the fifth. Using full sentences describe how to fully simplify Megan and Julie’s expressions. Describe if Megan and Julie started with equivalent expressions or if they started with expressions that are not equal.

OpenStudy (anonymous):

@SolomonZelman can you help me?

OpenStudy (solomonzelman):

Megan \(\LARGE\color{blue}{ \bf \frac{x^{1/3}}{x^{12}} }\) Julie \(\LARGE\color{blue}{ \bf \sqrt[32]{x\times x^{2}\times x^{5}} }\)

OpenStudy (solomonzelman):

Julie \(\LARGE\color{blue}{ \bf \sqrt[32]{x\times x^{2}\times x^{5}}= }\) \(\LARGE\color{blue}{ \bf \sqrt[32]{ x^{8}}= }\) \(\LARGE\color{blue}{ \bf x~^{8/32} =}\) \(\LARGE\color{blue}{ \bf x^{1/4} .}\)

OpenStudy (solomonzelman):

Megan \(\LARGE\color{blue}{ \bf \frac{ x^{1/3} }{ x^{12} } = }\) \(\LARGE\color{blue}{ \bf x^{1/3-12} }\) \(\LARGE\color{blue}{ \bf x~^{-11 ~~2/3} }\)

OpenStudy (solomonzelman):

The expressions are obviously not equivalent.

OpenStudy (solomonzelman):

obviously not equivalent, after simplifying of course.

OpenStudy (anonymous):

i see that, hmm thank you very much for working this out and showing each step i understand what you did to get there i appreciate your help :)